# Quantum Dynamics of Electron Scattering from Skyrmions

### v1

@@ II Theoretical Formulation
Figure 2: (a) The contour map of the potential components for 2D skyrmion based on Eq. (6). The color gradient indicates the potential strength. The skyrmion is centered at $(0.5L,0.5L)$ with a radius of $0.05L$ where $L$ is the length of the box. The off-diagonal terms are complex with $V^{\uparrow\downarrow}=V^{\downarrow\uparrow*}$(b) Potential components of 1D skyrmion centered at $0.5L$ with length $0.1L$. The spin texture of it is shown with arrows.
@@ III Numerical Recipe
Figure 3: The flow chart of the numerical methods described in Sec. II to solve the TDSE of the propagating electron in any arbitrary 2D potential. A finite difference method is followed, and a hybridized technique of operator-splitting with forward elimination and back substitution is employed. This approach reduces the computational complexity and enhances the numerical stability, providing an improvement over the conventional matrix inversion method [28].
@@ IV.1 Scattering from 1D skyrmion
Case 1- $\beta<1$ High kinetic energy: When the incident kinetic energy exceeds the critical energy scale $2J$, the spin-preserving component transmits through the potential region with minimal reflection (see Figs. 4-(c) and 4-(f) ). This minor reflection occurs due to the low-momentum modes present in the Gaussian tail. However, the spin-flip components, which are dynamically generated inside the skyrmion, despite having large kinetic energy, show higher reflection than transmission (see the inset plot in Figs. 4-(c) and 4-(f) ). This non-trivial phenomenon is due to the occurrence of a destructive interference between the forward and backward moving spin-flip components in the second half of the potential region. A detailed explanation of this destructive interference is provided in the supplementary material.
@@
Case 3- $\beta>1$ Low kinetic energy: As expected, there is a further decrease in transmission and an increase in reflection by lowering the kinetic energy. Most importantly, with increase in $\beta$ the transmission probability asymptotically decreases to a saturation. The underlying mechanism will be explained in the next paragraph. Furthermore, the metastable state is more prominent in this case. The double bump in the probability density of the spin-up electron and the kink in the spin-down electron (see Fig. 4-(o) ) emphasize our earlier explanation of the trapped spin-down state moving out as the spin-up state. This statement is corroborated by artificially disabling the $V^{\uparrow\downarrow}$ matrix element (not presented in the plots), which is responsible for down-to-up conversion, in which the spin-down state gets trapped inside the scatterer for all time. Further, in all cases of scattering, the spin-up electrons always move faster than their spin-down counterparts, which has been verified numerically by computing the expectation value of the kinetic energy as discussed in the supplementary material.
@@
In Fig. 5(a), we quantify the transmission and reflection, both for spin-preserved and spin-flip channels, as a function of $\beta$. The salient feature we get is that, in the limit $\beta\gg 1$, the reflection and transmission probabilities for the $\uparrow\uparrow$ channel saturate at finite values while the $\uparrow\downarrow$ probability vanishes. This finite saturation in the $\uparrow\uparrow$ channel is counterintuitive, as one expects the spin-up component to be almost completely reflected at large $\beta$. Two mechanisms could account for this behavior. Firstly, in the absence of a spin-flip potential, the problem reduces to the scattering of a spinless particle from a potential barrier, where the transmission probability asymptotically vanishes with a large barrier. Therefore, spin-flip is necessary to have a finite transmission probability. Secondly, there is a role of non-collinearity in spin texture. To disentangle it, in a hypothetical design, we suppress the non-collinearity while retaining the spin-flip coupling by adopting the following simplified model:
@@ IV.2 The 2D scattering
Building on the key features observed in the one-dimensional scattering scenario, in this section, we extend our analysis to scattering of electrons in a 2D skyrmion, the potential for which is shown in Fig. 2(a). While the 1D model offers meaningful insight into the core mechanisms of the scattering process, the transition to 2D brings additional insights. In the 1D case, all parts of the Gaussian wave packet are subject to the same potential profile, as shown in Fig. 2(a). However, in two dimensions, the presence of an azimuthal degree of freedom, $\phi$, introduces spatial anisotropy. Although different regions of the wave packet encounter qualitatively similar features of the potential, the effective height of the potential barrier varies with $\phi$, resulting in direction-dependent scattering outcomes that are absent in the 1D case. The results corresponding to scattering from 2D skyrmion for different energy regimes are shown in Fig. 6.
@@
Figure 6: Snapshot of time evolution of the electron wave function after scattering from a skyrmion for different $\beta=\frac{2J}{\text{KE}}$. This represents the 2D counterpart of 1D results shown in the third column of Fig. 4. The skyrmion region is marked by the black circle. The $45^{\circ}$ arrow shows the direction of initial propagation. The blue-magenta gradient represents the probability distribution of the spin-up particle, whereas the yellow-orange gradient represents the same for spin-down. The maximum of the color gradients is set at 2.0 for a clear visualization of the scattered state with lower probability.
@@
Irrespective of the value of $\beta$ and initial spin polarization, the large fraction of spin-preserving channels moves along the forward direction, and a small portion backscatters. However, the fraction of back scattering increases with an increase in $\beta$ as expected. This can be inferred from the higher number of wrinkles in $\uparrow\uparrow$$\downarrow\downarrow$ channel as we increase $\beta$ (see Fig. 6). We may note that some back scattering at low $\beta$ appears due to the infrared tails of the Gaussian distribution and is less prominent in the figure.
@@
Unlike the spin-preserving channels (see Fig. 6), the spin-flip channels are highly dependent on $\beta$ following Eq. (6). The scattering intensity of the $\downarrow\uparrow$ channel along the forward direction increases with $\beta$. However, it is not the case in the $\uparrow\downarrow$ channel. The reason is explained as follows. Even though the probability of $\uparrow\downarrow$ flipping is the same as in the case of $\downarrow\uparrow$ (see Eq.  (6)), the dynamically generated spin-down electrons inside the skyrmion via spin-flipping have to overcome the $2J$ barrier in order to propagate outside of the scatterer. Therefore, regardless of the fact that $\uparrow\downarrow$ conversion increases with $\beta$, due to a lack of sufficient kinetic energy, the spin-down states stay inside, except for some quantum tunneling, as can be seen in the $\uparrow\downarrow$ plot in the lower panel of Fig. 6. Nevertheless, these trapped spin-down states escape the barrier by back flipping to the spin-up state. This phenomenon can be observed from secondary waves whose elongated tails are present inside the skyrmion (see the $\uparrow\uparrow$ channels in Fig. 6). Furthermore, the presence of the angular phase factor $e^{\pm{i\phi}}$ introduces a spin to orbital momentum transfer in the spin-flip channel dynamically. This occurs to conserve the total angular momentum ($i\partial_{\alpha}+m\hat{s}_{z}$). We may note that real-time spin-angular momentum transfer has not been reported yet.
@@
In the scattered state, the expectation value of the kinetic energy of the spin-up electron is always higher than that of the spin-down state, as the former slides downhill and the latter has to climb uphill, as shown in Fig. 2, and the same has been validated numerically as discussed in the supplementary material.
@@ V.1 Summary
In this present work, the time evolution of a spin-polarized Gaussian wave packet of electrons has been studied by explicitly solving the TDSE numerically. The finite difference method is followed in our numerics up to first order in time, and an iterative propagation scheme has been implemented based on forward elimination and back substitution, along with the operator splitting technique. Using the designed numerical framework, we have studied the scattering dynamics of the electron wave packets interacting with chiral spin textures of skyrmions in both 1D and 2D geometries. Our simulations have revealed that the scattering process is insensitive to the chirality, $\gamma$. However, in agreement with the prior observations [24], the differential scattering cross section for skyrmion and anti-skyrmion exhibits a mirror symmetry along the wave packet’s propagation axis. A key advantage of our approach lies in its ability to capture the full spatio-temporal evolution of the wave packet, thereby providing physical insights into how the competition between the kinetic energy (KE) and skyrmion potential arising from the Hund’s coupling ($J$) between itinerant electrons and the local spin controls the spin-dependent scattering dynamics. Consequently, an important parameter $\beta(=\frac{2J}{KE})$ emerged to describe the scattering phenomena in this quantum system.
@@ Acknowledgements
Author contributions
All the authors together conceived the idea. HS developed the numerical methods. The results are analyzed by all the authors. The manuscript was prepared by HS, AM, and BRKN, and it was further revised by SS.
Competing interests
The authors declare no competing interests.

### v2

# arXiv:2509.19437v2[cond-mat.mes-hall] 02 Mar 2026

@@ II Theoretical Formulation
Figure 2: (a) The contour map of the potential components for 2D skyrmion based on Eq. (6). The color gradient indicates the potential strength. The skyrmion is centered at $(0.5L,0.5L)$ with a radius of $0.05L$ where $L$ is the length of the box. The off-diagonal terms are complex with $V^{\uparrow\downarrow}=V^{\downarrow\uparrow*}$. (b) A 3D schematic diagram for a 2D skyrmion. The marked rectangle indicates a 1D cross-section of the 2D skyrmion, which we defined as 1D skyrmion. (c) Potential components of this 1D skyrmion centered at $0.5L$ with length $0.1L$.
@@ III Numerical Recipe
Figure 3: The flow chart of the numerical methods described in Sec. II to solve the TDSE of the propagating electron in any arbitrary 2D potential. A finite difference method is followed, and a hybridized technique of operator-splitting with forward elimination and back substitution is employed. This approach reduces the computational complexity and enhances the numerical stability as explained in Appendix-C, which provides an improvement over the conventional matrix inversion method [10.1119/1.13811].
@@ IV.1 Scattering from 1D skyrmion
Case 1- $\beta<1$ High kinetic energy: When the incident kinetic energy exceeds the critical energy scale $2J$, the spin-preserving component transmits through the potential region with minimal reflection (see Figs. 4-(c) and 4-(f) ). This minor reflection occurs due to the low-momentum modes present in the Gaussian tail. However, the spin-flip components, which are dynamically generated inside the skyrmion, despite having large kinetic energy, show higher reflection than transmission (see the inset plot in Figs. 4-(c) and 4-(f) ). This non-trivial phenomenon is due to the occurrence of a destructive interference between the forward and backward moving spin-flip components in the second half of the potential region. A detailed explanation of this destructive interference is provided in the Appendix B material.
@@
Case 3- $\beta>1$ Low kinetic energy: As expected, there is a further decrease in transmission and an increase in reflection by lowering the kinetic energy. Most importantly, with an increase in $\beta$ the transmission probability asymptotically decreases to a saturation. The underlying mechanism will be explained in the next paragraph. Furthermore, the metastable state is more prominent in this case. The double bump in the probability density of the spin-up electron and the kink in the spin-down electron (see Fig. 4-(o) ) emphasize our earlier explanation of the trapped spin-down state moving out as the spin-up state. This statement is corroborated by artificially disabling the $V^{\uparrow\downarrow}$ matrix element (not presented in the plots), which is responsible for down-to-up conversion, in which the spin-down state gets trapped inside the scatterer for all time. Further, in all cases of scattering, the spin-up electrons always move faster than their spin-down counterparts, which has been verified numerically by computing the expectation value of the kinetic energy as discussed in the Appendix A material.
@@
In Fig. 5(a), we quantify the transmission and reflection, both for spin-preserved and spin-flip channels, as a function of $\beta$. The salient feature we get is that, in the limit $\beta\gg 1$, the reflection and transmission probabilities for the $\uparrow\uparrow$ channel (spin up electron scattered as spin up electron) saturate at finite values while the $\uparrow\downarrow$ (spin up electron scattered as spin down electron)probability vanishes. This finite saturation in the $\uparrow\uparrow$ channel is counterintuitive, as one expects the spin-up component to be almost completely reflected at large $\beta$. Two mechanisms could account for this behavior. Firstly, in the absence of a spin-flip potential, the problem reduces to the scattering of a spinless particle from a potential barrier, where the transmission probability asymptotically vanishes with a large barrier. Therefore, spin-flip is necessary to have a finite transmission probability. Secondly, there is a role of non-collinearity in spin texture. To disentangle it, in a hypothetical design, we suppress the non-collinearity while retaining the spin-flip coupling by adopting the following simplified model:
@@ IV.2 The 2D scattering
Building on the key features observed in the one-dimensional scattering scenario, in this section, we extend our analysis to scattering of electrons in a 2D skyrmion, the potential for which is shown in Fig. 2(a). While the 1D model offers meaningful insight into the core mechanisms of the scattering process, the transition to 2D brings additional insights. In the 1D case, all parts of the Gaussian wave packet are subject to the same potential profile, as shown in Fig. 2(c). However, in two dimensions, the presence of an azimuthal degree of freedom, $\phi$, introduces spatial anisotropy. Although different regions of the wave packet encounter qualitatively similar features of the potential, the effective height of the potential barrier varies with $\phi$, resulting in direction-dependent scattering outcomes that are absent in the 1D case. The results corresponding to scattering from 2D skyrmion for different energy regimes are shown in Fig. 6.
@@
Figure 6: Snapshot of time evolution of the electron wave function after scattering from a skyrmion for different $\beta=\frac{2J}{\text{KE}}$. The real-time dynamics is provided in supplementary material [2] This represents the 2D counterpart of 1D results shown in the third column of Fig. 4. The skyrmion region is marked by the black circle. The $45^{\circ}$ arrow shows the direction of initial propagation. The blue-magenta gradient represents the probability distribution of the spin-up particle, whereas the yellow-orange gradient represents the same for spin-down. The maximum of the color gradients is set at 2.0 for a clear visualization of the scattered state with lower probability.
@@
Irrespective of the value of $\beta$ and initial spin polarization, the large fraction of spin-preserving channels moves along the forward direction, and a small portion backscatters. However, the fraction of back scattering increases with an increase in $\beta$ as expected. This can be inferred from the higher number of wrinkles in $\uparrow\uparrow$$\downarrow\downarrow$(spin down electron scattered as spin down electron) channel as we increase $\beta$ (see Fig. 6). We may note that some back scattering at low $\beta$ appears due to the infrared tails of the Gaussian distribution and is less prominent in the figure.
@@
Unlike the spin-preserving channels (see Fig. 6), the spin-flip channels are highly dependent on $\beta$ following Eq. (6). The scattering intensity of the $\downarrow\uparrow$(spin down electron scattered as spin up electron) channel along the forward direction increases with $\beta$. However, it is not the case in the $\uparrow\downarrow$ channel. The reason is explained as follows. Even though the probability of $\uparrow\downarrow$ flipping is the same as in the case of $\downarrow\uparrow$ (see Eq.  (6)), the dynamically generated spin-down electrons inside the skyrmion via spin-flipping have to overcome the $2J$ barrier in order to propagate outside of the scatterer. Therefore, regardless of the fact that $\uparrow\downarrow$ conversion increases with $\beta$, due to a lack of sufficient kinetic energy, the spin-down states stay inside, except for some quantum tunneling, as can be seen in the $\uparrow\downarrow$ plot in the lower panel of Fig. 6. Nevertheless, these trapped spin-down states escape the barrier by back flipping to the spin-up state. This phenomenon can be observed from secondary waves whose elongated tails are present inside the skyrmion (see the $\uparrow\uparrow$ channels in Fig. 6). Furthermore, the presence of the angular phase factor $e^{\pm{i\phi}}$ introduces a spin to orbital momentum transfer in the spin-flip channel dynamically. This occurs to conserve the total angular momentum ($i\partial_{\alpha}+m\hat{s}_{z}$). We may note that real-time spin-angular momentum transfer has not been reported yet.
@@
In the scattered state, the expectation value of the kinetic energy of the spin-up electron is always higher than that of the spin-down state, as the former slides downhill and the latter has to climb uphill, as shown in Fig. 2, and the same has been validated numerically as discussed in the Appendix A material.
@@ V.1 Summary
In this present work, the time evolution of a spin-polarized Gaussian wave packet of electrons has been studied by explicitly solving the TDSE numerically. The finite difference method is followed in our numerics and an iterative propagation scheme has been implemented based on forward elimination and back substitution, along with the operator splitting technique. Our numerical scheme has second order convergence both in space and time as shown in Appendix C. Using the designed numerical framework, we have studied the scattering dynamics of the electron wave packets interacting with chiral spin textures of skyrmions in both 1D and 2D geometries. Our simulations have revealed that the scattering process is insensitive to the chirality, $\gamma$. However, in agreement with the prior observations [10], the differential scattering cross section for skyrmion and anti-skyrmion (skyrmion with opposite vorticity) exhibits a mirror symmetry along the wave packet’s propagation axis. A key advantage of our approach lies in its ability to capture the full spatio-temporal evolution of the wave packet, thereby providing physical insights into how the competition between the kinetic energy (KE) and skyrmion potential arising from the Hund’s coupling ($J$) between itinerant electrons and the local spin controls the spin-dependent scattering dynamics. Consequently, an important parameter $\beta(=\frac{2J}{KE})$ emerged to describe the scattering phenomena in this quantum system.
@@ Acknowledgments
Data Availibility
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
@@ Appendix B The near-zero transmission in the spin-flip channel
Figure 9: (a)The plot of the total probability as a function of time steps, and the inset plot shows the fluctuations. This indicates that our numerical scheme conserves the total probability to within an error of $10^{-6}$. (b) The log-log plot of the $E_{L_{2}}^{x}$ vs. spatial step size $\Delta x$. Magenta dots represent the computed errors, and the blue dashed line indicates the linear fit. (c) The log-log plot of the $E_{L_{2}}^{t}$ vs. temporal step size $\Delta t$. Magenta dots represent the computed errors, and the blue dashed line indicates the linear fit.
Appendix C Convergence test and error analysis
In order to comment on the accuracy of the numerical scheme, we plot the total probability as a function of time in Fig. 9(a), which is conserved to unity with accuracy up to the sixth decimal point. Note that this is a consequence of the splitting of the time evolution operator in Eq. 13 to preserve the unitarity. For further validation, we perform a convergence test on both the spatial and temporal grids by computing the $L_{2}$ (square integrable) error norm. Let us consider that $F(x_{i_{\rm{ref}}})$ is the reference function on the defined finest mesh for $X$ where $x_{i_{\rm{ref}}}\in X$ are the reference mesh points. Now if the coarse function $G$ available on the coarse mesh points $x_{i_{\rm{c}}}\in X$, then the $L_{2}$ error norm is defined as
$E_{L_{2}}^{x}=\sqrt{\sum_{i}\bigg|F(x_{i_{\rm{ref}}}=x_{i_{c}})-G(x_{i_{c}})\bigg|^{2}\Delta x}$ (54)
where $\Delta x$ is the uniform spatial step size on the coarse mesh. If
$E_{L_{2}}\sim(\Delta x)^{n}$ (55)
then, the spatial order of convergence is said to be $n$ [24, 8]. Below, we illustrate the convergence test for the case of the electron scattering from the 1D skyrmion. The analysis remains valid for the other cases presented here. The reference finest mesh is designed with total number of mesh points $N_{\rm{ref}}=6400$.
To determine the spatial order of convergence for our numerical scheme, we fix the time step size $\Delta t$ and the coarse meshes are taken to be $N_{c}=\frac{1}{\Delta x}=\{{200,400,800,1600\}}$. The log-log plot of $E_{L_{2}}$ vs. $\Delta x$ is plotted in Fig. 9(b). The slope is found to be $\sim 2.05$, determining the spatial order of convergence is $2$. This is because the spatial discretization of our differential operator
$\frac{\partial^{2}\psi}{\partial x^{2}}\approx\frac{\psi_{i+1}-2\psi_{i}+\psi_{i-1}}{\Delta x^{2}}+\mathcal{O}(\Delta x^{2})$ (56)
carries a leading order error term proportional to $\Delta x^{2}$.
Next, we compute the temporal order of convergence by fixing the spatial step size $\Delta x$ and varying the time step size $\Delta t$ by varying $\lambda$ in Eq. 16. We take the reference $\lambda_{\rm{ref}}=8$ and the coarse $\lambda_{c}=\{0.5,1,2,4\}$ and plotted a similar log-log plot of $E_{L_{2}}^{t}$ vs. $\Delta t$ in Fig. 9(c). The slope of the plot is found to be $\sim-2.12$, inferring a second-order convergence in time step, which is justified as follows. As we are using the trapezoidal integration rule where the local truncation error per step is $\sim\mathcal{O}(\Delta t^{3})$. However, the global error is
${\rm{global\,\,error}}$ $\sim N\times{\rm{local\,\,error}}$
$\sim\frac{1}{\Delta t}\times\mathcal{O}(\Delta t^{3})\sim\mathcal{O}(\Delta t^{2}).$ (57)
This is exactly what we observe in our plot.